formula for average intensity

Grade:11

9 Answers

PRAPULPODISHETTI
8 years ago
Average Intensity=1/2 BKWA2
B-BULK MODULUS.
 
KUNCHAM SAMPATH
8 years ago
1/2 BKWA2
B-BULK MODULUS.
Gavvala Ganesh
8 years ago
 
Average Intensity=1/2 BKWA2
B-BULK MODULUS.
KRANTHI
8 years ago
Average Intensity=1/2 BKWA2
B-BULK MODULUS.
RAKESH CHINDAM
8 years ago
Average Intensity=1/2 BKWA2
B-BULK MODULUS.
L GOUSE BASHA
8 years ago
AVERAGE INTENSITY ½ BKWA2
B IS BULK MODULUS
K IS KINETIC ENRGY
SEELA YUVARAJU
8 years ago
average intensity=1/2 BKWA2
B-balk modulus
vamshi
8 years ago
AVERAGE INTENSITY=½ BKWA2
  where , B is the BULK MODULUS
SAI SARDAR
8 years ago
AVERAGE INTENSITY  (I) =1/2 BKWA2 .
HERE,
B = BULK MODULUS.
A = AMPLTUDE.
 K = WAVE NUMBER
W = ANGULAR FREQUENCY.